OpenStudy (rossmeador):

Solve -log9 (x-8)=1 Solve -10log6 -9x=-40

OpenStudy (anonymous):

is that base 9?

OpenStudy (rossmeador):

Yes, base 9 and base 6

OpenStudy (anonymous):

x-8=1/9

OpenStudy (anonymous):

is that a -ve sign?

OpenStudy (rossmeador):

Negative? There are three negative terms, -10, -9x and -40 Could you also elaborate on how you come x-8=1/9 please?

OpenStudy (anonymous):

log9(x-8)=-1 x-8=9^-1

OpenStudy (anonymous):

log6(-9x)=4 -9x=6^4

OpenStudy (rossmeador):

Is x suppose to be isolated?

OpenStudy (anonymous):

wat do u mean by isolated?

OpenStudy (rossmeador):

Should I divide 6^4 by -9 so that x=-144